# Optimal Inapproximability of Promise Equations

### v2

@@ 5 Algebraic Approach to Max-PCSPs
$\left|\mathbb{E}_{f\sim\Omega^{\textnormal{Out}}}\mathbb{E}_{\mathbf{a}}\left[\mathrm{tr}\left(\left(\sum_{\tau\in\widehat{\mathscr{G}_{1}^{E}},\tau\neq 1}\sum_{s,t\in N_{\tau}}\dim_{\tau}\widehat{\mathcal{A}}(\tau_{s,t})\tau_{s,t}({\mathbf{a}})\right)\times\right.\right.\right.$
@@
For the tractability part, if $s/c\leq 1/|\textnormal{Im}(\varphi)|$ then the following algorithm works: Count the number equations that are unsatisfiable over $\mathscr{G}_{2}$. If the fraction of unsatisfiable equations is at most $1-c$, accept. In this case, the random assignment over $\textnormal{Im}(\varphi)$ satisfies at least a $c/|\textnormal{Im}(\varphi)|$-fraction of the equations in $\mathscr{G}_{2}$. Otherwise, reject.
For the hardness part, let $10<s\leq c<$ 1 be such that $s/c>1/|\textnormal{Im}(\varphi)|$, and let $\epsilon>0$ be a rational number such that $c<1-\epsilon<s|\textnormal{Im}(\varphi)|$. Then, the NP-hardness of $\operatorname{3-LIN}(\mathscr{G}_{1},\mathscr{G}_{2},\varphi,c,s)$ follows from the observation that
@@ Appendix B Proofs from Section 2
[…]
@@
[…]

### v3

@@ 5 Algebraic Approach to Max-PCSPs
[…]
@@
For the tractability part, if $s/c\leq 1/|\textnormal{Im}(\varphi)|$ then the following algorithm works: Count the number equations that are unsatisfiable over $\mathscr{G}_{2}$. If the fraction of unsatisfiable equations is at most 1c1-c, accept. In this case, the random assignment over $\textnormal{Im}(\varphi)$ satisfies at least a $c/|\textnormal{Im}(\varphi)|$-fraction of the equations in $\mathscr{G}_{2}$. Otherwise, reject.
For the hardness part, let $0<s\leq c<1$ be such that $s/c>1/|\textnormal{Im}(\varphi)|$, and let $\epsilon>0$ be a rational number such that $c<1-\epsilon<s|\textnormal{Im}(\varphi)|$. Then, the NP-hardness of $\operatorname{3-LIN}(\mathscr{G}_{1},\mathscr{G}_{2},\varphi,c,s)$ follows from the observation that
@@ Appendix B Proofs from Section 2
Let $\mathscr{H}\leq\mathscr{G}$ be groups, and let $\beta\in\widehat{\mathscr{H}}$. The representation induced by $\beta$ on $\mathscr{G}$, denoted $\alpha=\mathrm{Ind}_{\mathscr{H}}^{\mathscr{G}}\beta$, is defined as follows. Let $N_{\alpha}=N_{\beta}\times(\mathscr{H}\backslash\mathscr{G})$. Fix a set of representatives $\{g_{1},\dots g_{k}\}$ of $\mathscr{H}\backslash\mathscr{G}$. The matrix entries of $\alpha$ are given by
@@
Let $\alpha=\mathrm{Ind}_{\mathscr{H}}^{\mathscr{G}}1$. Then the index set of $\alpha$ is $\{1\}\times\mathscr{H}\backslash\mathscr{G}$, which can be naturally identified with $\mathscr{H}\backslash\mathscr{G}$. Under this identification, the matrix entries of $\alpha$ and $R_{\mathscr{G},\mathscr{H}}$ are the same, and the claim holds. Lemma 2.11 then follows from this claim and the following result. [^9] The formulation in [Ter99] is slightly different. To see that it is equivalent to ours, use [Ter99, Proposition 3.2] and the fact that we defined the inner product over $\mathcal{L}^{2}(\mathscr{G})$ with an additional normalising factor of $\frac{1}{|\mathscr{G}|}$.